A +100 μC charge is placed at the origin. A -50 μC is placed at x=2m and a +200 μC is placed at x= -4m.

What is the net electric force acting on the +100 μC charge?





1
Expert's answer
2021-04-05T11:08:04-0400

"F=F_1+F_2=kq_1q_2\/r_1^2+kq_3q_2\/r_2^2="



"=9\cdot10^9\cdot200\cdot10^{-6}\cdot100\cdot10^{-6}/4^2+"



"+9\cdot10^9\cdot50\cdot10^{-6}\cdot100\cdot10^{-6}/2^2=22.5\ (N)"